# Anti-rotors and algebra invariants

### v35

@@ 7.2. Two-dimensional algebras
Up to an isomorphism, there are only three two-dimensional algebras. For the algebra of complex numbers, $\mathbb{C}$, we can write its elements $s$ as $(x,y)$, with the product given by (x,y)(z,w)=(xzyw,yz+xw)(x,y)(z,w)=(xz-yw,yz+xw), and $\mathbf{1}=(1,0)$, so that $\|\mathbf{1}\|^{2}=1$. Also, $(x,y)^{-1}=(\frac{x}{x^{2}+y^{2}},\frac{-y}{x^{2}+y^{2}})$. We start out by writing $\mathfrak{u}_{\mathbb{C}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&\gamma\end{array}\right]$. We then obtain $\mathfrak{u}_{\mathbb{C}}s^{-1}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&\gamma\end{array}\right]\left(\begin{array}[]{ccc}x\\ y\end{array}\right)^{-1}=\left(\begin{array}[]{ccc}\frac{\alpha x-\beta y}{x^{2}+y^{2}}\\ \frac{\beta x-\gamma y}{x^{2}+y^{2}}\end{array}\right)$. As regards the right-hand-side of this equation, satisfaction of (3.1) requires that the partial derivative with respect to $x$ of the second component must equal the partial derivative with respect to $y$ of the first component. Consequently, $\mathfrak{u}_{\mathbb{C}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&-\alpha.\end{array}\right]$. Using Definition 4.1, a simple exercise in integration leads to $\log\mathfrak{u}^{d}_{\mathbb{C}}=\frac{\alpha}{2}\log(x^{2}+y^{2})+\beta{\rm arctan}(\frac{y}{x})$, relevant to the invariant reduced triples.
@@
$\log\mathfrak{u}^{d}_{\not{\mathbb{C}}}=\frac{\alpha}{2}\log[(x+y)(x-y)]+\frac{\beta}{2}\log\frac{x+y}{x-y}$. In fact, this can be simplified to $\log\mathfrak{u}^{d}_{\not{\mathbb{C}}}=\frac{\alpha+\beta}{2}\log(x+y)+\frac{\alpha-\beta}{2}\log(x-y)$. Of course, $\frac{\alpha+\beta}{2}$, $\frac{\alpha-\beta}{2}$ are themselves two independent parameters that we could name $\alpha^{\prime}$, $\beta^{\prime}$.
The final real two-dimensional unital algebra (up to an isomorphism) is the algebra of dual numbers $\mathbb{D}$. The elements are $(x,y)$ with product given by $(x,y)(z,w)=(xz,xw+yz)$, with $\mathbf{1}=(1,0)$ and $\|\mathbf{1}\|^{2}=1$. Also, $(x,y)^{-1}=(\frac{1}{x},-\frac{y}{x^{2}})$. We then find $\mathfrak{u}_{\mathbb{D}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&0\end{array}\right]$ and $\log\mathfrak{u}^{d}_{\mathbb{D}}=\alpha\log x+\beta\left(\frac{y}{x}\right)$.
Now we consider $\mathbb{R}\times\mathbb{R}$ (understood to be a direct product of rings), which is isomorphic to $\not{\mathbb{C}}$. We have elements as $(x,y)$, the product $(x,y)(z,w)=(xz,yw)$, and the multiplicative identity $\mathbf{1}=(1,1)$, so that $\|\mathbf{1}\|^{2}=2$. Also $(x,y)^{-1}=(\frac{1}{x},\frac{1}{y})$. We then find $\mathfrak{u}_{\mathbb{R}\times\mathbb{R}}=\left[\begin{array}[]{ccc}\alpha&0\\ 0&\gamma\end{array}\right]$ and $\log\mathfrak{u}^{d}_{\mathbb{R}\times\mathbb{R}}=\frac{\alpha}{2}\log x+\frac{\gamma}{2}\log y$.
Let us look further at the isomorphic algebras $\not{\mathbb{C}}$ and $\mathbb{R}\times\mathbb{R}$. It is clear that $\mathfrak{u}_{\not{\mathbb{C}}}$ and $\mathfrak{u}_{\mathbb{R}\times\mathbb{R}}$ are consistent based on the algebra isomorphism resulting from
$\left[\begin{array}[]{ccc}1&-1\\ 1&1\end{array}\right]\left(\begin{array}[]{ccc}x\\ y\end{array}\right)=\left(\begin{array}[]{ccc}x-y\\ x+y\end{array}\right)$, and $\log\mathfrak{u}^{d}_{\not{\mathbb{C}}}$, and $\log\mathfrak{u}^{d}_{\mathbb{R}\times\mathbb{R}}$ as calculated above are consistent (relevant to the invariant reduced triples).
@@
Table I. In this table, as well as in Table II and Table III, an expression with Greek letters is understood to imply the set whose members are all realizations of those parameters as real numbers. Each row indicates a representative of one of the three two-dimensional algebra isomorphism classes along with its left regular representation, followed by the associated anti-rotor and unital norm group ($\approx$” means “isomorphic to”). Note that $\not{\mathbb{C}}$ and $\mathbb{R}\times\mathbb{R}$ are isomorphic, their respective anti-rotors are compatible as regards Theorem 3.2 (via the algebra isomorphism $\left[\begin{array}[]{ccc}1&1\\ -1&1\end{array}\right]$), and their unital norm groups are compatible as regards the reduced triples (i.e., $(x+y)^{\frac{\alpha+\beta}{2}}(x-y)^{\frac{\alpha-\beta}{2}}$ versus $x^{\frac{\alpha}{2}}y^{\frac{\beta}{2}}$).
@@ 7.6.3. A five-dimensional algebra
8. Speculations regarding the complexity of this formalism
8.1. How “large” is the class of invariants implied by this approach, and how is the underlying formalism structured?
@@ 8. Speculations regarding the complexity of this formalism
While we have derived a mechanism that generates algebra isomorphism invariants based on selection of $f(s)=s^{-1}$ in Definition 3.1, there are infinitely many relevant possibilities for $f(s)$ that could substitute for $s^{-1}$. One might then wonder if the association of an algebra with its resulting anti-rotor “variants” (each resulting from a different choice of $f(s)$) could thereby be an endless source of algebra invariants. If so, it would be appropriate to wonder how one could bring order to a such a collection of processes that literally spew out invariants. Would there be different classes of functions that would generate different classes of invariants? Should one search for an organizing structure analogous to that gloriously provided by (co)homology theories? And so on.
@@
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. If in Definition 3.1 we replace $s^{-1}$ with $s^{j}$ for small $|j|$, $j\in\mathbb{Z}$, $j\neq 0,1$ (i.e., excluding the cases that are not dependent on the algebra product), it is quickly discovered that all resulting anti-rotor variants for the algebras in Table III are identical within each choice of algebra (where $n$ is small, if applicable).
@@
This result seems surprising. That is, let $\hat{s}$ be the left regular representation of an algebra with elements $s$, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. For any uncurling metric $L$ (i.e., such that $Ls^{-1}$ is irrotational), the behavior described in the prior paragraph implies that the various vector fields defined by $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls$ are simultaneously irrotational, at least for small $|j|$ and small $n$. Thus, repetitive applications of $L\hat{s}L^{-1}$ or $L\hat{s}^{-1}L^{-1}$ initially seeded with the irrotational vector field $Ls^{-1}$ always result in irrotational vector fields.
@@
One might incorrectly suspect that this follows because repeated applications of an operator that resulted in an irrotational vector field will continue to result in irrotational vector fields. However, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. In general, this does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational. Of course, an important distinction between $L\hat{s}L^{-1}$ and $F(x,y)$ is that the former represents an algebra and the latter does not. So, generalization of the phenomenon described in the second paragraph of this section, if valid, would seem to depend on the presence of an algebra. But how would that work?
8.2. Algebras, Calculus, 1-forms, and pullbacks
Proceeding heuristically, the following analogy might be considered. We often refer to “critical phenomena” occurring with smooth nonconstant and nonlinear entities. A simple example is that of a smooth nonconstant and nonlinear function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$. When the gradient of the function vanishes at a point $\mathbf{x^{\prime}}$, the latter is a critical point of $g$. On a domain that does not include zeros of the function, if a function is raised to a power such that the result remains nonconstant and nonlinear, then $\mathbf{x^{\prime}}$ remains a critical point, and no new critical points are introduced (as is clear from the chain rule).
Does something like this occur for vector fields and 1-forms? Normally, one cannot talk about raising these to a power - but such an operation does make sense in the context of an algebra with elements $s$ (as always, the standard basis is assumed). Thus, consider the nonconstant and nonlinear vector field $s^{j}$, $j\in\mathbb{Z}$, $j\neq 0,1$. If an alternative metric $\mu$ is introduced, then in the basis defined by the columns of $\mu$ the algebra elements are $\mu s$, and we have $(\mu s)^{j}=\mu s^{j}$, as well as $(\mu s^{j})^{r}=\mu s^{jr}$. Let us now introduce the function $f_{s^{j}}(\mu)\equiv\mu s^{j}$, which maps nonzero metrics to nonconstant and nonlinear vector fields. We might say that a “critical phenomenon” occurs at $\mu^{\prime}$ if, with respect to a suitable domain in $\mathbb{R}^{n}$, the 1-form dual to $\mu^{\prime}s^{j}$ is exact, i.e., if the exterior derivative of the corresponding 1-form vanishes (or in other words, if the curl of $\mu^{\prime}s^{j}$ vanishes there). Making an analogy with Calculus, we might then refer to $\mu^{\prime}$ as a critical point of $f_{s^{j}}(\mu)$. As with the behavior of critical phenomenon concerning the smooth nonconstant and nonlinear function $g$, expressed at the end of the prior paragraph, we might expect that the set of critical points of $f_{s^{j}}(\mu)=\mu s^{j}$ and $\left(f_{s^{j}}(\mu)\right)^{r}=\mu s^{jr}$ are the same, so long as $\left(f_{s^{j}}(\mu)\right)^{r}$ again maps metrics to vector fields that are nonconstant and nonlinear (which happens if $r$ is a nonzero integer). If so, that would explain the phenomenon described in the second paragraph of Section 8.1, and indicate that it generalizes.
The above analogy at least suggests the following approach. The chain rule in Calculus generalizes to the setting of the prior paragraph in the statement that “pullbacks commute with differential forms”, i.e., $d(F^{*}\omega)=F^{*}d\omega$. Thus, consider the assertion that for any $j\in\mathbb{Z}$, $j\neq 0,1$, there is a diffeomorphism $F_{j}$ on a relevant domain of $\mathbb{R}^{n}$ that results in a pullback between the two 1-forms $(\mu s^{j})\cdot\mathbf{d}s$ and $(\mu s^{-1})\cdot\mathbf{d}s$. If one could prove that assertion, it would explain the phenomenon of the second paragraph of Section 8.1. That is, we would have $(\mu s^{j})\cdot\mathbf{d}s=F_{j}^{*}\big{(}(\mu s^{-1})\cdot\mathbf{d}s\big{)}$, along with a similar equation expressing $\big{(}\mu s^{-1}\big{)}\cdot\mathbf{d}s$ as a pullback of $\big{(}\mu s^{j}\big{)}\cdot\mathbf{d}s$. Commutativity of a pullback and a differential form would then imply that the exterior derivative of $\mu^{\prime}s^{j}\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $\mu^{\prime}s^{-1}\cdot\mathbf{d}s$ vanishes.
On the other hand, proving that those pullbacks exist would involve a demonstration of the existence of solutions to large algebra-dependent collection of nonlinear partial differential equations.
8.3. Are all anti-rotor variants the same?
Leaving aside the speculations of Section 8.2, the limited computational evidence concerning the algebras in Table III cited in Section 8.1 does raise the possibility that, for a given algebra, identical anti-rotors result from replacing $s^{-1}$ in Definition 3.1 with $s^{j}$, for any $j\in\mathbb{Z}$ except for $j=0$ or $j=1$ (i.e., replacing $s^{-1}$ with any $s^{j}$ that actually depends on the algebra product). In view of the linearity of the exterior derivative, it would then similarly be the case that $f(s)$ given by different linear combinations of integer powers of $s$ will yield the same anti-rotor, so long as the linear combinations include one power that is not 0 or 1. Now suppose one deals with the left regular representation of an algebra. Recall that if a matrix substitutes for the variable in a Laurent series of a function, then the result converges if the matrix eigenvalues lie in the annulus of convergence of the Laurent series. With that in mind we can propose the following.
Conjecture.
Consider a function of a complex variable $f(z)$ that is holomorphic in an annulus centered at z=0, and its Laurent series there has at least one nonzero coefficient for a power of $z$ not equal to 0 or 1, and all coefficients in the series are real. Then for an algebra $A$ that is a subalgebra of $M_{n}(\mathbb{R})$, there is some domain $D_{A,f}$ in its vector space of elements such that the anti-rotor variant obtained by modifying Definition 3.1, where $f(s)$ replaces $s^{-1}$ and $D_{A,f}$ replaces the neighborhood of $\mathbf{1}$, yields the same result as the unaltered Definition 3.1.
We summarize as follows. In Definition 3.1 we can replace $s^{-1}$ with $f(s)$ for $f(z)$ a function that is holomorphic in an annulus $D_{A,f}$ centered at the origin, and replace the neighborhood of $\mathbf{1}$ with $D_{A,f}$. The purview of our formalism thus consists of all invariants resulting from each of the infinite number of resulting anti-rotor “variants”. On its face, this is a chaotic-appearing situation that one would immediately seek to tame in some possibly complicated manner. However, if the above Conjecture is proved to be true (or the phenomenon described in the second paragraph of Section 8.1 generalizes), one is entitled to conclude that an algebra is truly associated with a unique anti-rotor (i.e., all anti-rotor variants are actually the same), and one only need deal with Definition 3.1 as written. The potential complexity of the formalism would thereby be hugely reduced.
Whether of not the Conjecture is true, anti-rotors of isomorphic algebras are related according to Theorem 3.2, leading to recognition of the various invariants identified in this paper, to be added to an algebra’s dimension, number of subgroups, types of ideals, and invariants stemming from composition series and (co)homology.

### v36

@@ 7.2. Two-dimensional algebras
Up to an isomorphism, there are only three two-dimensional algebras. For the algebra of complex numbers, $\mathbb{C}$, we can write its elements $s$ as $(x,y)$, with the product given by (x,y)(z,w)=(xzyw,yz+xw)(x,y)(z,w)=(xz-yw,yz+xw), and $\mathbf{1}=(1,0)$, so that $\|\mathbf{1}\|^{2}=1$. Also, $(x,y)^{-1}=(\frac{x}{x^{2}+y^{2}},\frac{-y}{x^{2}+y^{2}})$. We start out by writing $\mathfrak{u}_{\mathbb{C}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&\gamma\end{array}\right]$. We then obtain $\mathfrak{u}_{\mathbb{C}}s^{-1}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&\gamma\end{array}\right]\left(\begin{array}[]{ccc}x\\ y\end{array}\right)^{-1}=\left(\begin{array}[]{ccc}\frac{\alpha x-\beta y}{x^{2}+y^{2}}\\ \frac{\beta x-\gamma y}{x^{2}+y^{2}}\end{array}\right)$. As regards the right-hand-side of this equation, satisfaction of (3.1) requires that the partial derivative with respect to $x$ of the second component must equal the partial derivative with respect to $y$ of the first component. Consequently, $\mathfrak{u}_{\mathbb{C}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&-\alpha.\end{array}\right]$. Using Definition 4.1, a simple exercise in integration leads to $\log\mathfrak{u}^{d}_{\mathbb{C}}=\frac{\alpha}{2}\log(x^{2}+y^{2})+\beta{\rm arctan}(\frac{y}{x})$. The invariant reduced triple (7.2) is then (0,0,1).
@@
$\log\mathfrak{u}^{d}_{\not{\mathbb{C}}}=\frac{\alpha}{2}\log[(x+y)(x-y)]+\frac{\beta}{2}\log\frac{x+y}{x-y}$. In fact, this can be simplified to $\log\mathfrak{u}^{d}_{\not{\mathbb{C}}}=\frac{\alpha+\beta}{2}\log(x+y)+\frac{\alpha-\beta}{2}\log(x-y)$. Of course, $\frac{\alpha+\beta}{2}$, $\frac{\alpha-\beta}{2}$ are themselves two independent parameters that we could name $\alpha^{\prime}$, $\beta^{\prime}$. Either way, the reduced triple (7.2) is (0,1,0).
The final real two-dimensional unital algebra (up to an isomorphism) is the algebra of dual numbers $\mathbb{D}$. The elements are $(x,y)$ with product given by $(x,y)(z,w)=(xz,xw+yz)$, with $\mathbf{1}=(1,0)$ and $\|\mathbf{1}\|^{2}=1$. Also, $(x,y)^{-1}=(\frac{1}{x},-\frac{y}{x^{2}})$. We then find $\mathfrak{u}_{\mathbb{D}}=\left[\begin{array}[]{ccc}\alpha&\beta\\ \beta&0\end{array}\right]$ and $\log\mathfrak{u}^{d}_{\mathbb{D}}=\alpha\log x+\beta\left(\frac{y}{x}\right)$. The reduced triple (7.2) is (1,0,0).
Now we consider $\mathbb{R}\times\mathbb{R}$ (understood to be a direct product of rings), which is isomorphic to $\not{\mathbb{C}}$. We have elements as $(x,y)$, the product $(x,y)(z,w)=(xz,yw)$, and the multiplicative identity $\mathbf{1}=(1,1)$, so that $\|\mathbf{1}\|^{2}=2$. Also $(x,y)^{-1}=(\frac{1}{x},\frac{1}{y})$. We then find $\mathfrak{u}_{\mathbb{R}\times\mathbb{R}}=\left[\begin{array}[]{ccc}\alpha&0\\ 0&\gamma\end{array}\right]$ and $\log\mathfrak{u}^{d}_{\mathbb{R}\times\mathbb{R}}=\frac{\alpha}{2}\log x+\frac{\gamma}{2}\log y$. The reduced triple (7.2) is then (0,1,0).
$\frac{1}{2}\left[\begin{array}[]{ccc}1&-1\\ 1&1\end{array}\right]$ provides the isomorphism between the algebras $\not{\mathbb{C}}$ and $\mathbb{R}\times\mathbb{R}$. It is easy to verify that their anti-rotors as identified above are consistent as regards to Theorem 3.2, and we have seen that the two algebras have the same reduced triple.
@@
Table I. In this table, as well as in Table II and Table III, an expression with Greek letters is understood to imply the set whose members are all realizations of those parameters as real numbers. Each row indicates a representative of one of the three two-dimensional algebra isomorphism classes along with its left regular representation, followed by the associated anti-rotor and unital norm group ($\approx$” means “isomorphic to”). Note that $\not{\mathbb{C}}$ and $\mathbb{R}\times\mathbb{R}$ are isomorphic, their respective anti-rotors are compatible as regards Theorem 3.2 (via the algebra isomorphism $\frac{1}{2}\left[\begin{array}[]{ccc}1&1\\ -1&1\end{array}\right]$), and their unital norm groups are compatible as regards the reduced triples (i.e., $(x+y)^{\frac{\alpha+\beta}{2}}(x-y)^{\frac{\alpha-\beta}{2}}$ versus $x^{\frac{\alpha}{2}}y^{\frac{\beta}{2}}$).
@@ 7.6.3. A five-dimensional algebra
8. How large is the class of invariants implied by this approach?
@@
While we have derived a mechanism that generates algebra isomorphism invariants based on selection of $f(s)=s^{-1}$ in Definition 3.1, there are infinitely many relevant possibilities for $f(s)$ that could substitute for $s^{-1}$. One might then wonder if the association of an algebra with its resulting anti-rotor “variants” (each resulting from a different choice of $f(s)$) could thereby be an endless source of algebra invariants. If so, it would be appropriate to wonder how one could bring order to a such a collection of processes that literally spew out invariants. Would there be different classes of functions that would generate different classes of invariants? And so on.
@@
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. If in Definition 3.1 we replace $s^{-1}$ with $s^{j}$ for small $|j|$, $j\in\mathbb{Z}$, $j\neq 0,1$ (i.e., excluding the cases that are not dependent on the algebra product), it is quickly discovered that all resulting anti-rotor variants for the algebras in Table III are identical within each choice of algebra. Extrapolating, if the above in fact holds for all $j\in\mathbb{Z}$, $j\neq 0,1$, linearity of the exterior derivative means it would hold for arbitrary linear combinations of the various $s^{j}$ - implying that with respect to suitable domains in $\mathbb{R}^{n}$ it would hold for functions expressible by infinite series in the powers of $s$.
However, the behavior cited in the first sentence of the prior paragraph was unanticipated and seems very surprising. There are several obvious avenues one might think would lead to an understanding of it.
@@
First, let $\hat{s}$ be the left regular representation of an algebra with elements $s$, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. For any uncurling metric $L$ (i.e., such that $Ls^{-1}$ is irrotational), extrapolation of the behavior described in the penultimate paragraph suggests that the various vector fields defined by $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls$ are simultaneously irrotational. That is, repetitive applications of $L\hat{s}L^{-1}$ or $L\hat{s}^{-1}L^{-1}$ initially seeded with the irrotational vector field $Ls^{-1}$ always result in irrotational vector fields.
@@
One might incorrectly suspect that this follows because repeated applications of an operator that resulted in an irrotational vector field will continue to result in irrotational vector fields. However, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. In general, this does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational. Of course, an important distinction between $L\hat{s}L^{-1}$ and $F(x,y)$ is that the former are elements of an algebra and the latter are not. So, generalization of the phenomenon described in the second paragraph of this section, if valid, would seem to depend on the presence of an algebra. But how would that work?
Another possibility for understanding the behavior cited in the second paragraph of this subsection is suggested by what happens in the case of algebra $M_{n}(\mathbb{R})$, with respect to which there is the well known Matrix Calculus expression,
(8.1) $\frac{\partial}{\partial s}\left(\text{Tr}(s^{j})\right)=j\left(s^{j-1}\right)^{T},$
where Tr indicates trace (in the Matrix Calculus literature, $s$ in (8.1) is typically written as $\mathbf{X}$). Using Jacobi’s formula, we have noted earlier that the transpose operation T is an uncurling metric for $M_{n}(\mathbb{R})$ in the exceptional $f(s)=s^{-1}$ case, and (8.1) indicates that T is indeed also an uncurling metric for $f(s)=s^{j}$ in general. So, according to the proof of Proposition 4.5, here we have an example of an algebra whose space of uncurling metrics (in this case, one-dimensional) is independent of the choice of $f(s)=s^{j}$, for $j\neq 0,1$. Now, if we deal with the left regular representation of an algebra, then we will be working with subalgebras of $M_{n}(\mathbb{R})$ (and their usually larger spaces of uncurling metrics). In the latter setting, perhaps there is there an analogy to be exploited as regards the suggestive (8.1). But how would that work?
As a third way to try to find an explanation for the behavior cited in the second paragraph of this subsection, suppose it can be shown that a pullback deriving from a diffeomorphism $\phi_{\mu,j}$ exists between the 1-forms $(\mu s^{2})\cdot\mathbf{d}s$ and $(\mu s^{j})\cdot\mathbf{d}s$ when either of them is exact - i.e., whenever $\mu$ is an uncurling metric for either $s^{2}$ or $s^{j}$, $j\neq 0,1$. That is, we would have $(\mu s^{j})\cdot\mathbf{d}s=\phi_{\mu,j}^{*}\big((\mu s^{2})\cdot\mathbf{d}s\big)$, along with a similar equation expressing $\big(\mu s^{2}\big)\cdot\mathbf{d}s$ as a pullback of $\big(\mu s^{j}\big)\cdot\mathbf{d}s$. Commutativity of a pullback and a differential form would then imply,
$d\left((\mu s^{j})\cdot\mathbf{d}s\right)=d\left(\phi_{\mu,j}^{*}(\mu s^{2})\cdot\mathbf{d}s\right)=\phi_{\mu,j}^{*}d\big((\mu s^{2})\cdot\mathbf{d}s\big).$
This would mean that the exterior derivative of $(\mu s^{j})\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $(\mu s^{2})\cdot\mathbf{d}s$ vanishes - and we would have our explanation.
However, proving that those pullbacks exist would involve a demonstration of the existence of solutions to large algebra-dependent collections of nonlinear partial differential equations. For individual cases of a particular algebra, a particular uncurling metric $\mu^{\prime}$, and a particular value of $j$, along with the (unwelcome) stipulation that the diffeomorphisms must have only rational function components, the problem might be addressed with known techniques of computational algebraic geometry. But it is difficult to overcome the impression that the general problem is intractable, i.e., that of establishing the existence of the requisite pullbacks for any algebra and any $j\neq 0,1$.
Thus, the problem of explaining the behavior cited in the second paragraph of this subsection appears daunting. So we are left with the following, which for convenience addresses the left regular representations of algebras.
Conjecture 8.1.
Suppose $f(z)$ and $g(z)$ are two functions of a complex variable that are holomorphic in some annulus around the origin, all coefficients of their respective Laurent series are real, and each of these series have at least one term $\alpha_{j}z^{j}$ where $\alpha_{j}\neq 0$ for $j\neq 0,1$. Let $q$-dimensional algebra $A$ be a subalgebra of $M_{n}(\mathbb{R})$. Consider a domain $D$ in $A$ where all members of $\{s:s\in D\}\cup\{s^{-1}:s\in D\}$ are such that with respect to their matrix format their eigenvalues lie in the above annulus. Then with respect to $D$, and any given metric $\mu\in\text{Sym}_{q}$, $\mu f(s)$ is irrotational if and only if $\mu g(s)$ is irrotational, where $f(s)$ and $g(s)$ are treated as $q$-dimensional vector fields.
In essence, recalling Section 2, the conjecture asserts that the kernel of $\not{\text{S}}_{f}$ is independent of $f$ over the class of functions that depend on the algebra product.
To summarize: On its face, it would seem very unlikely that Conjecture 8.1 is true. And yet, one can generate a great deal of computational evidence in its favor - beginning with the algebras listed in Table III. But either way, resolution of the conjecture will lead to an interesting result.
  • If Conjecture 8.1 is true, then every algebra is uniquely associated with an anti-rotor (all anti-rotor variants arising from different relevant choices of $f$ are actually the same). This would be a nice outcome for what initially seemed to be a disorderly situation. Moreover, the truth of Conjecture 8.1 might be a deep result, as it presently seems so non-intuitive.
  • If Conjecture 8.1 is false, the set of invariants implied by our program will be (perhaps much) greater than we presently suspect, and new organizing principles will likely need to be developed to bring order and understanding to the full set of invariants generated by this program.